# Mixed Density Diffuser: Efficient Planning with Non-uniform Temporal Resolution

### v1

@@ Abstract
Recent studies demonstrate that diffusion planners benefit from sparse-step planning over single-step planning. Training models to skip steps in their trajectories helps capture long-term dependencies without additional or memory computational cost. However, predicting excessively sparse plans degrades performance. We hypothesize this temporal density threshold is non-uniform across a temporal horizon and that certain parts of a planned trajectory should be more densely planned. We propose Mixed Density Diffuser (MDD), a diffusion planner where the densities throughout the horizon are tunable hyperparameters. MDD achieves a new SOTA across the Maze2D, Franka Kitchen, and Antmaze D4RL task domains.
@@ 1 Introduction
To overcome the limitations of both uniform-density and hierarchical planners, we introduce Mixed Density Diffuser (MDD). Our primary contributions are:
@@ 2 Diffusion Planning
Diffusion planners [13, 14, 15, 17] approach offline RL as a guided sampling problem. Given a current state $s_{t}$ or state-action pair $(s_{t},a_{t})$, diffusion planners predict an $H$ step trajectory, $\tau$, with some jump variable $K$ timestep between each planned step. The initial position $X_{t}$ is fixed to the current true observation. A generalized trajectory output where $X_{t}$ can either be $s_{t}$ or $(s_{t},a_{t})$ is,
$\mathbf{\tau}=\begin{bmatrix}X_{t}&X_{t+K}&X_{t+2K}&\dots&X_{t+HK}\end{bmatrix}$
@@
3 Mixed Density Diffuser Framework
While our approach can be flexibly integrated with many diffusion planner designs we leverage the DV framework, the product of a comprehensive hyperparameter sweep [17]. MDD uses a DiT denoising backbone[22], MCSS guidance, and a single flat planner predicting state-only trajectories paired and a diffusion-based inverse dynamics model.
We can abstract the output trajectory as,
$\mathbf{\tau}=\begin{bmatrix}x_{t}&x_{t+K_{1}}&x_{t+K_{1}+K_{2}}&\dots&x_{t+K_{1}+K_{2}+\dots+K_{H-1}}\end{bmatrix}$
where each $K_{i}$ is an independent, tunable hyperparameter. Adjusting these $K_{i}$ jump sizes allows precise control of temporal resolution. Tuning ranges of jump indices quickly narrowed our hyperparameter search and is easier to interpret.
@@ 5 Discussion
Our framework adjustments only relate to the temporal density of the output trajectory states. Similarly to the Diffusion Veteran baselines, we use the same temporal density configuration for each task in a task domain demonstrating some hyperparameter robustness. Specifically, we use a denser-to-sparser approach. However, specific density configurations perform much higher on specific subtasks. A sparse-to-dense approach actually surpassed every model we compared in the Antmaze-Diverse task–scoring $84.0$. We theorize that adding higher temporal resolution to a latter part of the output trajectory gave the model more computational resources and training feedback for some long-term dependency, but leave further investigation for future work. Although MDD is outperformed slightly by $1$ to $2$ models on $4$ subtasks, the base DV framework was outperformed by an even larger margin on half of MDD’s underperforming subtasks. This suggests a single mixed density does not always generalize well to every subtask and that MDD inherits some of DV’s shortcomings. Particular, MDD and DV achieved very high performance on the Large and Medium Maze2D tasks, but were surpassed by two baselines on the Umaze task. The Umaze environment is smaller and simpler than the Large and Medium environments suggesting that approaches like HD and DQL have superior short-term planning but lack the long-horizon foresight to achieve a higher mean score on the Maze2D tasks. In both the Kitchen and Maze2D tasks, we compare two variations of the Diffusion Veteran framework to test whether MDD’s performance increase came from simply increasing the sparsity and temporal horizon of generated DV trajectories. We adopt the notation DV-J where J is the adjusted jump size. Surprisingly, we find that increasing DV’s jump size actually significantly increases it’s Kitchen performance despite DV having already been tuned in a hyperparameter sweep of over $6,000$ models [17]. MDD still outperforms these DV-J variants on average.

### v2

# Mixed-Density Diffuser: Efficient Planning with Non-uniform Temporal Resolution

@@ Abstract
Recent studies demonstrate that diffusion planners benefit from sparse-step planning over single-step planning. Training models to skip steps in their trajectories helps capture long-term dependencies without additional or memory computational cost. However, predicting excessively sparse plans degrades performance. We hypothesize this temporal density threshold is non-uniform across a temporal horizon and that certain parts of a planned trajectory should be more densely planned. We propose Mixed-Density Diffuser (MDD), a diffusion planner where the densities throughout the horizon are tunable hyperparameters. MDD achieves a new SOTA across the Maze2D, Franka Kitchen, and Antmaze D4RL task domains.
@@ 1 Introduction
To overcome the limitations of both uniform-density and hierarchical planners, we introduce Mixed-Density Diffuser (MDD). Our primary contributions are:
@@ 2 Diffusion Planning
Diffusion planners [13, 14, 15, 17] approach offline RL as a guided sampling problem. Given a current state $s_{t}$ or state-action pair $(s_{t},a_{t})$, diffusion planners predict an $H$ step trajectory, $\boldsymbol{\tau}$, with some jump variable $K$ timestep between each planned step. The initial position $X_{t}$ is fixed to the current true observation. A generalized trajectory output where $X_{t}$ can either be $s_{t}$ or $(s_{t},a_{t})$ is,
$\boldsymbol{\tau}=\begin{bmatrix}X_{t}&X_{t+K}&X_{t+2K}&\dots&X_{t+HK}\end{bmatrix}$
@@
3 Mixed-Density Diffuser Framework
While our approach can be flexibly integrated with many diffusion planner designs, we build on the DV framework, the result of a comprehensive hyperparameter sweep [17]. MDD employs a DiT denoising backbone [22], MCSS guidance, and a single flat planner that predicts state-only trajectories, paired with a diffusion-based inverse dynamics model.
We can abstract the output trajectory as:
$\boldsymbol{\tau}=\begin{bmatrix}x_{t}&x_{t+K_{1}}&x_{t+K_{1}+K_{2}}&\dots&x_{t+K_{1}+K_{2}+\dots+K_{H-1}}\end{bmatrix}$
where each $K_{i}$ is an independent, tunable hyperparameter. Adjusting these $K_{i}$ jump sizes allows precise control over temporal resolution. For example, our kitchen model denoises an output trajectory of $H=32$ entries, where $K_{i}=4$ for $i=1$ through $i=10$ and $K_{i}=6$ for $i=11$ through $31$. Manually tuning ranges of jump indices quickly narrowed our hyperparameter search and improved interpretability.
@@ 5 Discussion
Our framework adjustments only relate to the temporal density of the output trajectory states. Similarly to the Diffusion Veteran baselines, we use the same temporal density configuration for each task in a task domain demonstrating some hyperparameter robustness. Specifically, we use a denser-to-sparser approach. However, specific density configurations perform much higher on specific subtasks. A sparse-to-dense approach actually surpassed every model we compared in the Antmaze-Diverse task–scoring $84.0$. We theorize that adding higher temporal resolution to a latter part of the output trajectory gave the model more computational resources and training feedback for some long-term dependency, but leave further investigation for future work. Although MDD is outperformed slightly by $1$ to $2$ models on $4$ subtasks, the base DV framework was outperformed by an even larger margin on half of MDD’s underperforming subtasks. This suggests a single mixed-density does not always generalize well to every subtask and that MDD inherits some of DV’s shortcomings. Particular, MDD and DV achieved very high performance on the Large and Medium Maze2D tasks, but were surpassed by two baselines on the Umaze task. The Umaze environment is smaller and simpler than the Large and Medium environments suggesting that approaches like HD and DQL have superior short-term planning but lack the long-horizon foresight to achieve a higher mean score on the Maze2D tasks. In both the Kitchen and Maze2D tasks, we compare two variations of the Diffusion Veteran framework to test whether MDD’s performance increase came from simply increasing the sparsity and temporal horizon of generated DV trajectories. We adopt the notation DV-J where J is the adjusted jump size. Surprisingly, we find that increasing DV’s jump size actually significantly increases it’s Kitchen performance despite DV having already been tuned in a hyperparameter sweep of over $6,000$ models [17]. MDD still outperforms these DV-J variants on average.